Solve equations by factoring
Solve the equation.
1430x = 130x^2

Answer by ewatrrr(24785)


Hi,
Solving equation by factoring
1430x = 130x^2
130x^2 - 1430x = 0
130x(x - 11) = 0
130x = 0
x = 0
(x - 11) = 0
x = 11